# The Avengers: A Simple Recipe for Uniting Smaller Language Models to Challenge Proprietary Giants

### v2

@@ Abstract
As proprietary giants increasingly dominate the race for ever-larger language models, a pressing question arises for the open-source community: can smaller models remain competitive across a broad range of tasks? In this paper, we present the Avengers—a simple recipe that effectively leverages the collective intelligence of open-source, smaller language models. Our framework is built upon four lightweight operations: (i) embedding: encode queries using a text embedding model; (ii) clustering: group queries based on their semantic similarity; (iii) scoring: scores each model’s performance within each cluster; and (iv) voting: improve outputs via repeated sampling and voting. At inference time, each query is embedded and assigned to its nearest cluster. The top-performing model(s) within that cluster are selected to generate the response using the Self-Consistency or its multi-model variant. Remarkably, with 10 open-source models ($\sim$7B parameters each), the Avengers collectively outperforms GPT-4.1 on nine out of 15 datasets (spanning mathematics, code, logic, knowledge, and affective tasks). In particular, it surpasses GPT-4.1 on mathematics tasks by 18.21% and on code tasks by 7.46%. Furthermore, the Avengers delivers superior out-of-distribution generalization, and remains robust across various embedding models, clustering algorithms, ensemble strategies, and values of its sole parameter—the number of clusters. We have open-sourced the code on GitHub: https://github.com/ZhangYiqun018/Avengers
@@ 4.1 Experimental Setup
We select baseline methods from two distinct technical directions (mixture- and router- based) to comprehensively evaluate the effectiveness of the Avengers. Additionally, to provide intuitive performance benchmarks, we include a proprietary model baseline, gpt-4.1-2025-04-14 (GPT-4.1). Specifically, we consider the following baseline methods:

### v3

@@ Abstract
Proprietary giants are increasingly dominating the race for ever-larger language models. Can open-source, smaller models remain competitive across a broad range of tasks? In this paper, we present the Avengers—a simple recipe that leverages the collective intelligence of these smaller models. The Avengers builds upon four lightweight operations: (i) embedding: encode queries using a text embedding model; (ii) clustering: group queries based on their semantic similarity; (iii) scoring: scores each model’s performance within each cluster; and (iv) voting: improve outputs via repeated sampling and voting. At inference time, each query is embedded and assigned to its nearest cluster. The top-performing model(s) within that cluster are selected to generate the response with repeated sampling. Remarkably, with 10 open-source models ($\sim$7B parameters each), the Avengers surpasses GPT-4o, 4.1, and 4.5 in average performance across 15 diverse datasets spanning mathematics, coding, logical reasoning, general knowledge, and affective tasks. In particular, it surpasses GPT-4.1 on mathematics tasks by 18.21% and on code tasks by 7.46%. Furthermore, the Avengers delivers superior out-of-distribution generalization, and remains robust across various embedding models, clustering algorithms, ensemble strategies, and values of its sole parameter—the number of clusters. We have open-sourced the code on GitHub: https://github.com/ZhangYiqun018/Avengers
Figure 1: Comparison of proprietary models with our method (the Avengers) across six representative benchmarks. Full results on all 15 benchmarks are reported in Table 2.
@@ 4.1 Experimental Setup
We select baseline methods from two distinct technical directions (mixture- and router- based) to comprehensively evaluate the effectiveness of the Avengers. Additionally, to provide intuitive performance benchmarks, we include three proprietary models, gpt-4o-2024-08-06 (GPT-4o), gpt-4.5-preview (GPT-4.5-preview) and gpt-4.1-2025-04-14 (GPT-4.1). Specifically, we consider the following baseline methods: